OpenStudy (calculusxy):

MEDAL for help in Algebra 1! Simplify the following...

OpenStudy (calculusxy):

\[\large \frac{x+8}{27x^3 - 9x^2} - 3\]

OpenStudy (sweetburger):

first step is getting the terms to have a common denominator

OpenStudy (calculusxy):

The common denominator is 27x^3 - 9x^2

OpenStudy (amilapsn):

yep

OpenStudy (calculusxy):

So should I factor it?

OpenStudy (sweetburger):

no not yet

OpenStudy (calculusxy):

Ok.

OpenStudy (sweetburger):

your going to be expanding so might as well not factor it

OpenStudy (calculusxy):

So the first fraction you leave the numerator as it is. Second fraction, 3(27x^3 -9x^2) = 81x^3 - 27x^2

OpenStudy (sweetburger):

yep looks good. dont forget the negative sign though

OpenStudy (calculusxy):

oh yes so then it would be -81x^3 + 27x^2

OpenStudy (sweetburger):

yep and now you can combine the numerator over the single denominator

OpenStudy (calculusxy):

\(\huge \frac{x + 8 -81x^2 + 27x^2 }{27x^3 - 9x^2}\)

OpenStudy (sweetburger):

looks good except the term -81x^2 i think should be -81x^3

OpenStudy (calculusxy):

Oh yes. But don't I need to factor the denominator? It doesn't look too simplified

OpenStudy (calculusxy):

@sweetburger @agent0smith @ganeshie8 @KamiBug

OpenStudy (sweetburger):

if you factor the denominator what are you going to do with it? Unless you determine that you could possibly factor the numerator and cancel something out?